I need help with this problem. thank you so much
4/8 divided by 4/3

Answer by stanbon(75887)

4/8 divided by 4/3
------------
(4/8) / (4/3)
= (1/2) / (4/3)
Invert the denominator and multiply to get:
= (1/2)*(3/4)
= 3/8
=============
Cheers,
Stan H.